# Optimizing Sparse SYK

### v2

@@ 2.5 Commutation index and Lovász theta function
We follow closely the formulation of [4], though similar ideas have been advanced in works on studying uncertainty relations [17, 50], random Hamiltonians [5], and shadow tomography [33]. The commutation index is defined as follows:
@@
It is easy to see that the commutation index bounds the energy variance of any given state. In turn, the commutation index is known to be bounded by the Lovász theta function of the commutation graph of $\mathcal{S}$. This fact was proven in [17, Equation 5], [26], and [50, Proposition 3]. We define the Lovász theta function as follows, noting that other equivalent formulations exist as well.
@@
Lemma 2.11 (Bound on commutation index [17, 26, 50]).
@@
Lemma 2.12 (Bound on commutation index of SYK Hamiltonian [4]).
@@ 3 Maximum eigenvalue of sparse SYK Hamiltonians
SYK has an expected maximum eigenvalue of $\Theta(\sqrt{n})$ [26]. From Theorem 6 of [5], the expected maximum eigenvalue of $H_{SSYK}$ matches the standard SYK expected maximum eigenvalue when $n\rightarrow\infty$ under the condition that the disorder $\alpha_{I}=p^{-1/2}X_{I}J_{I}$ obeys the following conditions:
@@
Lemma 3.3 (Gaussian concentration inequality for Lipschitz functions; Theorem 2.26 of [46]).
@@
where the last inequality is from Lemma˜2.12 and holds for all $m$. As this Lipschitz bound holds for the energy over any state, it also holds for the supremum over all states which is the Lipschitz constant $L$ of $\lambda_{\mathrm{max}}(H_{SSYK}(p))$ is $O(n^{-1}p^{-1/2})$ (see, e.g., Proposition 8 of [5]). This gives:
@@
In Appendix˜A, we show an improvement can be made to the probability in which the upper bound holds when $2<a<3$ compared to Eq.˜34.
@@
again for $n\geq n_{0}$. Finally, we have
$\operatorname*{\mathbb{P}}\left[|\lambda_{\mathrm{max}}(H_{SSYK})-\operatorname*{\mathbb{E}}\lambda_{\mathrm{max}}(H_{SYK})\right|\geq t\sqrt{n})]$ $\leq\operatorname*{\mathbb{P}}\left[\left|\lambda_{\mathrm{max}}(H_{SSYK})-\operatorname*{\mathbb{E}}\lambda_{\mathrm{max}}(H_{SSYK})\right|\geq(t-t^{\prime})\sqrt{n}\right]$
$\leq 2\exp\left(-\Omega\left(\frac{(t-t^{\prime})^{2}n^{3}p}{2}\right)\right)$
@@
In the case of $a=3$, i.e. $p=\Theta(1/n^{3})$, we note how the existing algorithms proposed by [28, 29] outputting Gaussian states for sparse SYK when $p=\Theta(1/n^{3})$ achieve $\Theta(\sqrt{n})$. This shows that with probability exponentially close to 1, the maximum eigenvalue of the sparse SYK model is $\Theta(\sqrt{n})$ when $p=\Theta(1/n^{3})$, with the Gaussian state output of the above algorithms as a witness. However the expected maximum eigenvalue may differ from the dense SYK model by a constant factor, as opposed to the scaling in Lemma˜3.2.
@@ 4.1 Gaussian state energy upper bound
We begin by giving a detailed exposition of the results on SYK given in [25, Section VI]. We then extend the method to achieve results for sparse SYK.
@@
We have $\left\|\Gamma\right\|_{\mathrm{F}}=\sqrt{\operatorname*{\mathrm{Tr}}(\Gamma\Gamma^{\dagger})}=\sqrt{\operatorname*{\mathrm{Tr}}(\Gamma\Gamma^{\top})}=\sqrt{\operatorname*{\mathrm{Tr}}(\mathbb{I})}=\sqrt{2n}$ since $\Gamma$ is a real matrix and $\Gamma\Gamma^{\top}=\mathbb{I}$ for all pure fermionic Gaussian states. For Gaussian states that are not pure $\left\|\Gamma\right\|_{\mathrm{F}}\leq\sqrt{2n}$ by the same argument. Additionally, $J^{mat}_{p,4-p}$ is a $(2n)^{p}\times(2n)^{4-p}$ random matrix with i.i.d. standard Gaussian entries in its nonzero indices. Hastings bounds the operator norm of the random Gaussian matrices (e.g., using Theorem 4.4.3 of [45]), giving $\left\|J^{mat}_{0,4}\right\|_{\mathrm{op}}\leq O(1)$, $\left\|J^{mat}_{4,0}\right\|_{\mathrm{op}}\leq O(1)$, $\left\|J^{mat}_{1,3}\right\|_{\mathrm{op}}\leq O(1/\sqrt{n})$, $\left\|J^{mat}_{3,1}\right\|_{\mathrm{op}}\leq O(1/\sqrt{n})$, and $\left\|J^{mat}_{2,2}\right\|_{\mathrm{op}}\leq O(1/n)$, all with probability exponentially close to 1. This shows that
$\beta_{\sigma}\leq O(\sqrt{n})^{N_{L}+N_{R}}\cdot O(1)^{k_{0,4}+k_{4,0}}\cdot O(1/\sqrt{n})^{k_{1,3}+k_{3,1}}\cdot O(1/n)^{k_{2,2}}=O(n)^{k-k_{2,2}-k_{1,3}/2-k_{3,1}/2}O(1)^{k}.$ (41)
Hastings and Ryan O’Donnell show that there always exists some labeling of $\Gamma$ vectors as left and right such that $k_{2,2}=k$, guaranteeing that $\beta_{\sigma}\leq O(1)^{k}$ for any pairing $\pi$ [25]. There are at most $(4k-1)!!\leq(4k)^{2k}$ distinct pairings, and summing over all of them gives $\operatorname*{\mathrm{Tr}}(H^{k}\rho)\leq(4k)^{2k}\cdot O(1)^{k}=O(k)^{2k}$.
@@
We note that the random variable $p^{-1/2}X_{ij}J_{ij}$ has subgaussian norm $O(p^{-1/2})$, so a simple application of the operator norm bounds used by Hastings (e.g., Theorem 4.4.3 of [45]) would imply that $\left\|J^{X}_{2,2}\right\|_{\mathrm{op}}\leq 1/\Omega(np^{1/2})$ with probability $\geq 1-\exp(-\Omega(n))$. However, this gives an extra $\Omega(p^{-1/2})$ factor which we wish to avoid in the large $p$ regime. We instead use the following operator norm tail-bound:
Proposition 4.3 (Corollary 3.9 of [7]).
@@
From [25] (and as described above), we have $\lambda_{\mathrm{Gauss}}(H^{k}_{SSYK})$ is upper-bounded by $O(k)^{2k}\cdot O\left(n\left\|J^{X}_{2,2}\right\|_{\mathrm{op}}\right)^{k}$, where $J^{X}_{2,2}$ is a $4n^{2}\times 4n^{2}$ symmetric matrix with nonzero entries drawn from $\binom{2n}{4}^{-1/2}p^{-1/2}X_{ij}J_{ij}$ where $\{X_{ij}\}_{i\leq j}$ are i.i.d. Bernoulli random variables with $\operatorname*{\mathbb{P}}[X_{ij}=1]=p$ and $\{J_{ij}\}_{i\leq j}$ are standard i.i.d. Gaussian random variables. We will bound $\left\|J^{X}_{2,2}\right\|_{\mathrm{op}}$ with probability $1-\mathrm{poly}(n)$ over the draw of $H_{SSYK}$.
@@
Due to Hastings [25], we also get the following as a corollary from Lemma˜4.2 (see Appendix˜B):
@@ 4.2 Gaussian state energy lower bound
Considering the range for $p=\Theta(1/n^{a})$, where $2.5\leq a<3$, we note that the algorithm from [29] outputs a Gaussian state with approximation ratio $\Omega(1/k)$, where $k$ is the maximum degree of the interaction graph, i.e. the maximum number of terms any single Majorana operator is in. From Theorem˜3.1, this corresponds to energy $\Omega(\sqrt{n}/k)$. With sparsification parameter $p$, the degree of each vertex is $\text{Binomial}\left(O(n^{3}),p\right)$, which is upper-bounded by $O(n^{3}p)$ with probability exponentially close to 1 by standard concentration bounds (see Subsection˜2.1). Taking a union bound over all $O(n^{4})$ vertices, we have $k\leq O(n^{3}p)$ with probability exponentially close to 1. Conditioned on this event, the energy of the Gaussian state outputted by the algorithm from [29] is $\Omega\left(\frac{\sqrt{n}}{n^{3}p}\right)=\Omega\left(\frac{1}{pn^{2.5}}\right)$.
Our argument for the $\Omega(1)$ energy lower bound is similar to Remark 7.14 in the arXiv version of [26], though we reproduce the full details here for completeness. Note that the Gaussian state depends on the specific SSYK Hamiltonian. Let $\bm{g}^{0}$ be the $2n\times 2n$ block diagonal matrix:
@@
where $\bm{1}\left\{\cdot\right\}$ denotes the indicator function. The bottom-right $n\times n$ block of $\bm{g}^{1}$ defines a Wigner matrix with entries with variance $\Theta\left(np\right)$. Standard results in random matrix theory (see [6] or Theorem 2.1.22 of [3]) then imply that $\left\|\bm{g}^{1}\right\|_{\mathrm{op}}=\Theta\left(n\sqrt{p}\right)$ w.h.p.
@@ 4.3 Low circuit complexity energy upper bound
With probability exponentially close to 1 over the draw of the Hamiltonian, any collection of states $S$ containing a state which achieves energy $t$ on the sparse SYK model with sparsity parameter $p$ must have at least $|S|\geq\exp(\Omega(pn^{2}t^{2}))$ distinct states.
@@
Figure 4: The empirical scaling of the Lovász number of sparsified commutation graphs with respect to the sparsification parameter $p$ conforms well to a scaling of $\Theta(\sqrt{p})$. Here, we plot a comparison of the Lovász number of sparsified commutation graphs of the SYK model with their fit to a square root power law $c_{1}\sqrt{p}+c_{2}$ where $c_{1}$ and $c_{2}$ are the fit parameters. Error bars are equal to the standard deviation over values drawn from $10$ random sparsifications of the initial graph for each data point.
@@
With probability exponentially close to 1 over the draw of the Hamiltonian, any collection of states $S$ containing a state which achieves energy $t$ with high probability on the sparse SYK model with sparsity parameter $p$ must have at least $|S|\geq\exp(\Omega(\sqrt{p}n^{2}t^{2}))$ distinct states.
@@ 5 Optimal energy quantum algorithm for sparse SYK Hamiltonians
where the $X_{S,j}$ are i.i.d. Bernoulli random variables with $\operatorname*{\mathbb{P}}(X_{S,j}=1)=p$, $J_{S,j}$ are i.i.d. standard Gaussian random variables, $\phi_{S}\triangleq\prod_{i\in S}\phi_{i}$ for the ordered set $S$, and $n_{1},n_{2}=\Theta(n)$. We can see that this directly generalizes the original two-color (dense) SYK-4 model [26] by setting $p=1$. Note that we will write $H^{(2)}_{SSYK}$ and $\tau_{j}$ where the dependence on $p$ is implicitly assumed.
@@ 5.2 Hastings–O’Donnell algorithm applied to sparse SYK
The Hastings–O’Donnell Algorithm [26] is a quantum algorithm for finding a constant approximation to the optimal energy state for the two-color SYK Hamiltonian. We show that this algorithm also works on the sparse two-color SYK Hamiltonian when $p\geq\Omega(\log n/n)$.
@@
In [26] it was shown that this state achieves with high probability a constant approximation ratio for the dense ($p=1$) two-color SYK model for some constant $\theta$. This algorithm is summarized as Algorithm˜1 and a visual diagram is shown in Fig.˜5.
@@
For any $p\geq\Omega(\log n/n)$, there exists constants $\theta,C\in\mathbb{R}$ such that $\operatorname*{\mathrm{Tr}}(\rho_{\theta}H_{SSYK}^{(2)})\geq C\sqrt{n}$ with probability at least $1-e^{-\Omega(n)}$ over the draw of $H_{SSYK}^{(2)}$.
@@
Following the high-level idea from [26], we calculate the energy of $\rho_{\theta}$ on $H_{SSYK}^{(2)}$.
$\operatorname*{\mathrm{Tr}}\left(\rho_{\theta}H_{SSYK}^{(2)}\right)=\operatorname*{\mathrm{Tr}}\left(e^{-\theta\zeta}\rho_{0}e^{+\theta\zeta}H_{SSYK}^{(2)}\right)=\operatorname*{\mathrm{Tr}}\left(\rho_{0}e^{+\theta\zeta}H_{SSYK}^{(2)}e^{-\theta\zeta}\right)=\operatorname*{\mathrm{Tr}}\left(\rho_{0}\mathrm{Ad}_{e^{+\theta\zeta}}(H_{SSYK}^{(2)})\right)$ (66)
@@
$\frac{2}{\sqrt{n_{2}}}\cdot 2^{-(n_{2}+n_{1}/2)}\sum_{j=1}^{n_{2}}\operatorname*{\mathrm{Tr}}(\tau_{j}^{2})$ $=\frac{2}{\sqrt{n_{2}}}\cdot 2^{-(n_{2}+n_{1}/2)}\sum_{j=1}^{n_{2}}\binom{n_{1}}{3}^{-1}p^{-1}\operatorname*{\mathrm{Tr}}\left(\left(\sum_{\begin{subarray}{c}S\subseteq[n_{1}]\\ |S|=3\end{subarray}}X_{S}J_{S,j}\phi_{S}\right)^{2}\right)$
@@
Conditioned on the $X_{S}$, $\sum_{S}X_{S}J_{S,j}^{2}$ is a chi-squared random variable with degrees of freedom $d$, where $d\leq pn_{1}^{3}$ with high probability over the draw of the Binomial random variables $X_{S}$. The Laurent–Massart [37] upper-bound for chi-squared random variables gives:
@@
We define $J_{S,j}$ as a mean-zero Gaussian with variance 1 if $S$ has all distinct elements, and variance 0 otherwise. A direct calculation shown in [28] gives the value of $\zeta,[\zeta,H_{SSYK}^{(2)}]]$ as
$\begin{split}A\triangleq\frac{1}{\sqrt{n_{2}}}\binom{n_{1}}{3}^{-3/2}p^{-3/2}\sum_{j,k,\ell=1}^{n_{2}}\sum_{\begin{subarray}{c}S,S^{\prime},S^{\prime\prime}\subset[n_{1}]\\ |S|,|S^{\prime}|,|S^{\prime\prime}|=3\end{subarray}}f(S,S^{\prime},S^{\prime\prime},j,k,\ell)(X_{S,j}J_{S,j})(X_{S^{\prime},k}J_{S^{\prime},k})&(J_{S^{\prime\prime},\ell}X_{S^{\prime\prime},\ell})\\ \times\Big(\phi_{S}\sigma_{j}\phi_{S^{\prime}}\sigma_{k}\phi_{S^{\prime\prime}}\chi_{\ell}\Big)\end{split}$ (72)
@@
$A(1)\triangleq\frac{1}{\sqrt{n_{2}}}\binom{n_{1}}{3}^{-3/2}p^{-3/2}\sum_{j,k,\ell=1}^{n_{2}}\sum_{\begin{subarray}{c}S,S^{\prime},S^{\prime\prime}\subset[n_{1}]\\ |S|,|S^{\prime}|,|S^{\prime\prime}|=3\end{subarray}}f(S,S^{\prime},S^{\prime\prime},j,k,\ell)(X_{S,j}J_{S,j})(X_{S^{\prime},k}J_{S^{\prime},k})(X_{S^{\prime\prime},\ell}J_{S^{\prime\prime},\ell}),$ (73)
@@
We prove that Lemma˜5.7 implies Lemma˜5.6, following the moment method techniques outlined in [26, 28]. By Jensen’s inequality we have

### v3

@@ 2.5 Commutation index and Lovász theta function
We follow closely the formulation of [undefg], though similar ideas have been advanced in works on studying uncertainty relations [undefag, undefah], random Hamiltonians [undefai], and shadow tomography [undefaj]. The commutation index is defined as follows:
@@
It is easy to see that the commutation index bounds the energy variance of any given state. In turn, the commutation index is known to be bounded by the Lovász theta function of the commutation graph of $\mathcal{S}$. This fact was proven in [undefag, Equation 5], [undeff], and [undefah, Proposition 3]. We define the Lovász theta function as follows, noting that other equivalent formulations exist as well.
@@
Lemma 2.11 (Bound on commutation index [undefag, undeff, undefah]).
@@
Lemma 2.12 (Bound on commutation index of SYK Hamiltonian [undefg]).
@@ 3 Maximum eigenvalue of sparse SYK Hamiltonians
SYK has an expected maximum eigenvalue of $\Theta(\sqrt{n})$ [undeff]. From Theorem 6 of [undefai], the expected maximum eigenvalue of $H_{SSYK}$ matches the standard SYK expected maximum eigenvalue when $n\rightarrow\infty$ under the condition that the disorder $\alpha_{I}=p^{-1/2}X_{I}J_{I}$ obeys the following conditions:
@@
Lemma 3.3 (Gaussian concentration inequality for Lipschitz functions; Theorem 2.26 of [undefak]).
@@
where the last inequality is from Lemma˜2.12 and holds for all $m$. As this Lipschitz bound holds for the energy over any state, it also holds for the supremum over all states which is the Lipschitz constant $L$ of $\lambda_{\mathrm{max}}(H_{SSYK}(p))$ is $O(n^{-1}p^{-1/2})$ (see, e.g., Proposition 8 of [undefai]). This gives:
@@
In Appendix˜A, we show an improvement can be made to the probability in which the upper bound holds when $2<a<3$ compared to Eq.˜35.
@@
again for $n\geq n_{0}$. Finally, we have $\operatorname*{\mathbb{P}}\left[|\lambda_{\mathrm{max}}(H_{SSYK})-\operatorname*{\mathbb{E}}\lambda_{\mathrm{max}}(H_{SYK})\right|\geq t\sqrt{n})]\leq\operatorname*{\mathbb{P}}\left[\left|\lambda_{\mathrm{max}}(H_{SSYK})-\operatorname*{\mathbb{E}}\lambda_{\mathrm{max}}(H_{SSYK})\right|\geq(t-t^{\prime})\sqrt{n}\right]$ and
$\operatorname*{\mathbb{P}}\left[\left|\lambda_{\mathrm{max}}(H_{SSYK})-\operatorname*{\mathbb{E}}\lambda_{\mathrm{max}}(H_{SSYK})\right|\geq(t-t^{\prime})\sqrt{n}\right]$ $\leq 2\exp\left(-\Omega\left(\frac{(t-t^{\prime})^{2}n^{3}p}{2}\right)\right)$
@@
In the case of $a=3$, i.e. $p=\Theta(1/n^{3})$, we note how the existing algorithms proposed by [undefs, undeft] outputting Gaussian states for sparse SYK when $p=\Theta(1/n^{3})$ achieve $\Theta(\sqrt{n})$. This shows that with probability exponentially close to 1, the maximum eigenvalue of the sparse SYK model is $\Theta(\sqrt{n})$ when $p=\Theta(1/n^{3})$, with the Gaussian state output of the above algorithms as a witness. However the expected maximum eigenvalue may differ from the dense SYK model by a constant factor, as opposed to the scaling in Lemma˜3.2.
@@ 4.1 Gaussian state energy upper bound
We begin by giving a detailed exposition of the results on SYK given in [undefw, Section VI]. We then extend the method to achieve results for sparse SYK.
@@
We have $\left\|\Gamma\right\|_{\mathrm{F}}=\sqrt{\operatorname*{\mathrm{Tr}}(\Gamma\Gamma^{\dagger})}=\sqrt{\operatorname*{\mathrm{Tr}}(\Gamma\Gamma^{\top})}=\sqrt{\operatorname*{\mathrm{Tr}}(\mathbb{I})}=\sqrt{2n}$ since $\Gamma$ is a real matrix and $\Gamma\Gamma^{\top}=\mathbb{I}$ for all pure fermionic Gaussian states. For Gaussian states that are not pure $\left\|\Gamma\right\|_{\mathrm{F}}\leq\sqrt{2n}$ by the same argument. Additionally, $J^{mat}_{p,4-p}$ is a $(2n)^{p}\times(2n)^{4-p}$ random matrix with i.i.d. standard Gaussian entries in its nonzero indices. Hastings bounds the operator norm of the random Gaussian matrices (e.g., using Theorem 4.4.3 of [undefal]), giving $\left\|J^{mat}_{0,4}\right\|_{\mathrm{op}}\leq O(1)$, $\left\|J^{mat}_{4,0}\right\|_{\mathrm{op}}\leq O(1)$, $\left\|J^{mat}_{1,3}\right\|_{\mathrm{op}}\leq O(1/\sqrt{n})$, $\left\|J^{mat}_{3,1}\right\|_{\mathrm{op}}\leq O(1/\sqrt{n})$, and $\left\|J^{mat}_{2,2}\right\|_{\mathrm{op}}\leq O(1/n)$, all with probability exponentially close to 1. This shows that
$\beta_{\sigma}$ $\leq O(\sqrt{n})^{N_{L}+N_{R}}\cdot O(1)^{k_{0,4}+k_{4,0}}\cdot O(1/\sqrt{n})^{k_{1,3}+k_{3,1}}\cdot O(1/n)^{k_{2,2}}$
$\leq O(n)^{k-k_{2,2}-k_{1,3}/2-k_{3,1}/2}O(1)^{k}.$
Hastings and Ryan O’Donnell show that there always exists some labeling of $\Gamma$ vectors as left and right such that $k_{2,2}=k$, guaranteeing that $\beta_{\sigma}\leq O(1)^{k}$ for any pairing $\pi$ [undefw]. There are at most $(4k-1)!!\leq(4k)^{2k}$ distinct pairings, and summing over all of them gives $\operatorname*{\mathrm{Tr}}(H^{k}\rho)\leq(4k)^{2k}\cdot O(1)^{k}=O(k)^{2k}$.
@@
We note that the random variable $p^{-1/2}X_{ij}J_{ij}$ has subgaussian norm $O(p^{-1/2})$, so a simple application of the operator norm bounds used by Hastings (e.g., Theorem 4.4.3 of [undefal]) would imply that $\left\|J^{X}_{2,2}\right\|_{\mathrm{op}}\leq 1/\Omega(np^{1/2})$ with probability $\geq 1-\exp(-\Omega(n))$. However, this gives an extra $\Omega(p^{-1/2})$ factor which we wish to avoid in the large $p$ regime. We instead use the following operator norm tail-bound:
Proposition 4.3 (Corollary 3.9 of [undefam]).
@@
From [undefw] (and as described above), we have $\lambda_{\mathrm{Gauss}}(H^{k}_{SSYK})$ is upper-bounded by $O(k)^{2k}\cdot O\left(n\left\|J^{X}_{2,2}\right\|_{\mathrm{op}}\right)^{k}$, where $J^{X}_{2,2}$ is a $4n^{2}\times 4n^{2}$ symmetric matrix with nonzero entries drawn from $\binom{2n}{4}^{-1/2}p^{-1/2}X_{ij}J_{ij}$ where $\{X_{ij}\}_{i\leq j}$ are i.i.d. Bernoulli random variables with $\operatorname*{\mathbb{P}}[X_{ij}=1]=p$ and $\{J_{ij}\}_{i\leq j}$ are standard i.i.d. Gaussian random variables. We will bound $\left\|J^{X}_{2,2}\right\|_{\mathrm{op}}$ with probability $1-\mathrm{poly}(n)$ over the draw of $H_{SSYK}$.
@@
Due to Hastings [undefw], we also get the following as a corollary from Lemma˜4.2 (see Appendix˜B):
@@ 4.2 Gaussian state energy lower bound
Considering the range for $p=\Theta(1/n^{a})$, where $2.5\leq a<3$, we note that the algorithm from [undeft] outputs a Gaussian state with approximation ratio $\Omega(1/k)$, where $k$ is the maximum degree of the interaction graph, i.e. the maximum number of terms any single Majorana operator is in. From Theorem˜3.1, this corresponds to energy $\Omega(\sqrt{n}/k)$. With sparsification parameter $p$, the degree of each vertex is $\text{Binomial}\left(O(n^{3}),p\right)$, which is upper-bounded by $O(n^{3}p)$ with probability exponentially close to 1 by standard concentration bounds (see Subsection˜2.1). Taking a union bound over all $O(n^{4})$ vertices, we have $k\leq O(n^{3}p)$ with probability exponentially close to 1. Conditioned on this event, the energy of the Gaussian state outputted by the algorithm from [undeft] is $\Omega\left(\frac{\sqrt{n}}{n^{3}p}\right)=\Omega\left(\frac{1}{pn^{2.5}}\right)$.
Our argument for the $\Omega(1)$ energy lower bound is similar to Remark 7.14 in the arXiv version of [undeff], though we reproduce the full details here for completeness. Note that the Gaussian state depends on the specific SSYK Hamiltonian. Let $\bm{g}^{0}$ be the $2n\times 2n$ block diagonal matrix:
@@
where $\bm{1}\left\{\cdot\right\}$ denotes the indicator function. The bottom-right $n\times n$ block of $\bm{g}^{1}$ defines a Wigner matrix with entries with variance $\Theta\left(np\right)$. Standard results in random matrix theory (see [undefan] or Theorem 2.1.22 of [undefao]) then imply that $\left\|\bm{g}^{1}\right\|_{\mathrm{op}}=\Theta\left(n\sqrt{p}\right)$ w.h.p.
@@ 4.3 Low circuit complexity energy upper bound
With probability exponentially close to 1 over the draw of the Hamiltonian, any disorder-independent set of states $S$ containing a state which achieves energy $t$ on the sparse SYK model with sparsity parameter $p$ must have at least $|S|\geq\exp(\Omega(pn^{2}t^{2}))$ distinct states.
@@
Figure 4: The empirical scaling of the Lovász number of sparsified commutation graphs with respect to the sparsification parameter $p$ conforms well to a scaling of $\Theta(\sqrt{p})$. Here, we plot a comparison of the Lovász number of sparsified commutation graphs of the SYK model with their fit to a square root power law $c_{1}\sqrt{p}+c_{2}$ where $c_{1}$ and $c_{2}$ are the fit parameters. Error bars are equal to the standard deviation over values drawn from 1010 random sparsifications of the initial graph for each data point.
@@
With probability exponentially close to 1 over the draw of the Hamiltonian, any disorder-independent set of states $S$ containing a state which achieves energy $t$ with high probability on the sparse SYK model with sparsity parameter $p$ must have at least $|S|\geq\exp(\Omega(\sqrt{p}n^{2}t^{2}))$ distinct states.
@@ 5 Optimal energy quantum algorithm for sparse SYK Hamiltonians
where the $X_{S,j}$ are i.i.d. Bernoulli random variables with $\operatorname*{\mathbb{P}}(X_{S,j}=1)=p$, $J_{S,j}$ are i.i.d. standard Gaussian random variables, $\phi_{S}\triangleq\prod_{i\in S}\phi_{i}$ for the ordered set $S$, and $n_{1},n_{2}=\Theta(n)$. We can see that this directly generalizes the original two-color (dense) SYK-4 model [undeff] by setting $p=1$. Note that we will write $H^{(2)}_{SSYK}$ and $\tau_{j}$ where the dependence on $p$ is implicitly assumed.
@@ 5.2 Hastings–O’Donnell algorithm applied to sparse SYK
The Hastings–O’Donnell Algorithm [undeff] is a quantum algorithm for finding a constant approximation to the optimal energy state for the two-color SYK Hamiltonian. We show that this algorithm also works on the sparse two-color SYK Hamiltonian when $p\geq\Omega(\log n/n)$.
@@
In [undeff] it was shown that this state achieves with high probability a constant approximation ratio for the dense ($p=1$) two-color SYK model for some constant $\theta$. This algorithm is summarized as Algorithm˜1 and a visual diagram is shown in Fig.˜5.
@@
For any $p\geq\Omega(\log n/n)$, there exist positive constants $\theta,C\in\mathbb{R}$ such that $\operatorname*{\mathrm{Tr}}(\rho_{\theta}H_{SSYK}^{(2)})\geq C\sqrt{n}$ with probability at least $1-e^{-\Omega(n)}$ over the draw of $H_{SSYK}^{(2)}$.
@@
Following the high-level idea from [undeff], we calculate the energy of $\rho_{\theta}$ on $H_{SSYK}^{(2)}$.
$\operatorname*{\mathrm{Tr}}\left(\rho_{\theta}H_{SSYK}^{(2)}\right)=\operatorname*{\mathrm{Tr}}\left(e^{-\theta\zeta}\rho_{0}e^{+\theta\zeta}H_{SSYK}^{(2)}\right)=\operatorname*{\mathrm{Tr}}\left(\rho_{0}\mathrm{Ad}_{e^{+\theta\zeta}}(H_{SSYK}^{(2)})\right)$ (66)
@@
$\frac{2}{\sqrt{n_{2}}}\cdot 2^{-(n_{2}+n_{1}/2)}\sum_{j=1}^{n_{2}}\operatorname*{\mathrm{Tr}}(\tau_{j}^{2})$ $=\frac{2}{\sqrt{n_{2}}}\cdot 2^{-(n_{2}+n_{1}/2)}\sum_{j=1}^{n_{2}}\binom{n_{1}}{3}^{-1}p^{-1}\operatorname*{\mathrm{Tr}}\left((\sum_{\begin{subarray}{c}S\subseteq[n_{1}]\\ |S|=3\end{subarray}}X_{S}J_{S,j}\phi_{S})^{2}\right)$
@@
Conditioned on the $X_{S}$, $\sum_{S}X_{S}J_{S,j}^{2}$ is a chi-squared random variable with degrees of freedom $d$, where $d\leq pn_{1}^{3}$ with high probability over the draw of the Binomial random variables $X_{S}$. The Laurent–Massart [undefap] upper-bound for chi-squared random variables gives:
@@
We define $J_{S,j}$ as a mean-zero Gaussian with variance 1 if $S$ has all distinct elements, and variance 0 otherwise. A direct calculation shown in [undefs] gives the value of $\zeta,[\zeta,H_{SSYK}^{(2)}]]$ as
$\begin{split}A\triangleq\frac{1}{\sqrt{n_{2}}}\binom{n_{1}}{3}^{-3/2}p^{-3/2}\sum_{j,k,\ell=1}^{n_{2}}\sum_{\begin{subarray}{c}S,S^{\prime},S^{\prime\prime}\subset[n_{1}]\\ |S|,|S^{\prime}|,|S^{\prime\prime}|=3\end{subarray}}&f(S,S^{\prime},S^{\prime\prime},j,k,\ell)(X_{S,j}J_{S,j})(X_{S^{\prime},k}J_{S^{\prime},k})\\ &\times(J_{S^{\prime\prime},\ell}X_{S^{\prime\prime},\ell})\Big(\phi_{S}\sigma_{j}\phi_{S^{\prime}}\sigma_{k}\phi_{S^{\prime\prime}}\chi_{\ell}\Big)\end{split}$ (72)
@@
$A(1)\triangleq\frac{1}{\sqrt{n_{2}}}\binom{n_{1}}{3}^{-3/2}p^{-3/2}\sum_{j,k,\ell=1}^{n_{2}}\sum_{\begin{subarray}{c}S,S^{\prime},S^{\prime\prime}\subset[n_{1}]\\ |S|,|S^{\prime}|,|S^{\prime\prime}|=3\end{subarray}}$ $f(S,S^{\prime},S^{\prime\prime},j,k,\ell)(X_{S,j}J_{S,j})$
$\times(X_{S^{\prime},k}J_{S^{\prime},k})(X_{S^{\prime\prime},\ell}J_{S^{\prime\prime},\ell}),$
@@
We prove that Lemma˜5.7 implies Lemma˜5.6, following the moment method techniques outlined in [undeff, undefs]. By Jensen’s inequality we have